OpenStudy (anonymous):

Assume the y varies inversely with x and y = -3 when x = 4. Write the indirect variation equation for the relationship. Leave no spaces.

OpenStudy (aroub):

yx=k -3(4)=k -12=k

OpenStudy (anonymous):

so, it would be written as k=-12 ?

OpenStudy (lalaly):

y=-12/x

OpenStudy (lalaly):

or yx=-12

OpenStudy (aroub):

Yes